# BARI-TH/25-772

### v1

@@ 5 Conclusions
The search for BSM effects is currently facing the evidence that deviations from SM predictions exist but they are small. For example, the flavour anomalies are below the required level of significance to claim the evidence of new physics. Being confident that NP exists, we have pursued the aim of justifying the absence of large deviations. We have scrutinised channels where anomalies have been detected together with processes strictly forbidden in SM, namely lepton flavour violating modes, and the correlations among them.
In the proposed simple extension of the SM, the ABCD model, the new gauge boson $Z^{\prime}$ has flavour conserving and flavour violating generation dependent couplings to fermions. The model is based on a special solution of the gauge anomaly cancellation equations, which relates the quark and lepton couplings to $Z^{\prime}$. The flavour observables in the two sectors are therefore intertwined, producing nontrivial correlations.
As noticed at the formulation of the model [1], quark and lepton sectors mutually act to prevent large deviations from SM. This is a welcome feature in view of the small discrepancies with respect to SM displayed by measurements of the flavour observables. Moreover, the possibility of having lepton flavour violation implies that modes, such as those induced by $b\to s\ell_{1}^{-}\ell_{2}^{+}$ with $\ell_{1}\neq\ell_{2}$, can occur and are correlated to LFV leptonic modes, such as $\mu^{-}\to e^{-}\gamma$ and $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, and to the LFC modes $b\to s\ell^{-}\ell^{+}$ as well. We summarize below our main findings and indicate the corresponding plots.
  • We have shown that the NP contribution to the Wilson coefficients ${\rm Re}(C_{9}^{\rm NP})_{\ell\ell}$ and ${\rm Re}(C_{10}^{\rm NP})_{\ell\ell}$, $\ell=\mu,\,\tau$, is $\simeq 10\%$ of the corresponding SM coefficients. The analogous LFV coefficients ${\rm Re}(C_{9}^{\rm NP})_{\mu\tau}$ and ${\rm Re}(C_{10}^{\rm NP})_{\mu\tau}$ are approximately of the same size as ${\rm Re}(C_{9(10)}^{\rm NP})_{\mu\mu}$, as shown in Figs. 4-6.
  • We have found that LFV modes could reach detectable rates, probably accessible in the near future, see Figs. 14-16.
  • The role of quark and lepton observables to mutually constrain each other is visible in the results in Table 1 and in Figs. 17-18. The experimental upper limit for ${\cal B}(\mu\to e\gamma)$ efficiently constrains the $B$ and $B_{s}$ LFV decays.
  • The branching ratio of the LFV $\tau\to 3\mu$ mode can reach values close to the present upper bound, as shown in Fig. 19.
All findings are good news for the NP searches at the facilities planned within the future strategy for particle physics.
@@ Acknowledgements
F.D.F. thanks the coauthors of [1]. We thank A.J. Buras for very useful discussions. This study has been carried out within the INFN project (Iniziativa Specifica) SPIF. The research has been partly funded by the European Union – Next Generation EU through the research Grant No. P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Universitá e Ricerca (MUR).
@@ Appendix A $B_{q}\to V$ local form factors
$-\left.(\epsilon^{*}\cdot q)\frac{2m_{V}}{q^{2}}\big{(}A_{3}(q^{2})-A_{0}(q^{2})\big{)}q_{\mu}\right]\,\,,\,\,\,\hskip 14.22636pt$
@@
Table 2: Parameters of the $B_{(s)}\to V$ form factors in Eq. (A.5) determined in [60].
[…]
$+\Big{[}\epsilon^{*}_{\mu}(m_{B_{q}}^{2}-m^{2}_{V})-(\epsilon^{*}\cdot q)(p+p^{\prime})_{\mu}\Big{]}\;T_{2}(q^{2})\hskip 19.91684pt$
$+(\epsilon^{*}\cdot q)\left[q_{\mu}-\frac{q^{2}}{m_{B_{q}}^{2}-m^{2}_{V}}(p+p^{\prime})_{\mu}\right]\;T_{3}(q^{2})\,,\,\,\,\,\,\,$
@@
In our analysis we use the determination in [60] based on light-cone QCD sum rules (see also [80]). Defining the ”conformal” variable
@@
with parameters $\alpha_{k}^{i}$ and $m_{R,i}$. The value of $m_{R,i}$ is close to the mass of the lightest t-channel resonance involved in the various terms in (A.1), (A.2). The parameters are determined for $V,\,A_{1},\,A_{0},\,A_{12}$ and $T_{1},\,T_{2}$, $T_{23}$ in [60], defining
@@
To make easier to reproduce our results, we quote the values of the parameters in Table 2, while their correlation matrix is in [60].
@@ Appendix B Angular coefficient functions for $B\to K^{*}(K\pi)\ell_{1}^{-}\ell_{2}^{+}$
The coefficient functions $I_{i}^{(a)}$, for $i=1,\dots 9$ and $a=s,\,c$ in (3.1.2), can be written in terms of the transversity amplitudes
@@
with $\lambda_{H}=\lambda(q^{2},m_{B}^{2},m_{K^{*}}^{2})$ and $\lambda_{L}=\lambda(q^{2},m_{1}^{2},m_{2}^{2})$.
The expressions of the coefficient functions are:
@@ Appendix C Parametrization of the CKM and PMNS matrices
$V_{us}$ and $\gamma$ are set to $V_{us}=0.2253$ and $\gamma=64.6^{\circ}$. $|V_{cb}|$ and $|V_{ub}|$ are varied in the ranges spanned by the exclusive-inclusive measurements (69), with $|V_{cb}|_{exc}=(39.1\pm 0.5)\times 10^{-3}$, $|V_{cb}|_{inc}=(42.19\pm 0.78)\times 10^{-3}$, $|V_{ub}|_{exc}=(3.51\pm 0.12)\times 10^{-3}$, $|V_{ub}|_{inc}=(4.19\pm 0.17)\times 10^{-3}$ [5].
@@
where $x_{ij}=\sin\theta_{ij}$ and $k_{ij}=\cos\theta_{ij}$. A global fit of the parameters in (C.3) gives values which depend on the assumed hierarchy among the neutrino masses [81]: for normal ordering
@@
The global fit in [82] gives consistent results. In the numerical analysis we use the parameters for normal ordering (C.4).

### v2

@@ 5 Conclusions
The search for BSM is facing the evidence that tensions with the SM predictions are spotted, but they are small. For example, the flavour anomalies are below the level of significance to claim the evidence of new physics. Being confident on NP, we have attempted to justify the absence of large deviations: We have scrutinised channels where anomalies have been detected together with processes strictly forbidden in SM, the charged lepton flavour violating modes, and the correlations among them. In a simple extension of the SM, the ABCD model, the new gauge boson $Z^{\prime}$ has flavour conserving and flavour violating generation-dependent couplings to fermions. The model is based on a special solution of the gauge anomaly cancellation equations which relates the quark and lepton couplings to $Z^{\prime}$. The flavour observables in the two sectors are intertwined, producing nontrivial correlations.
As noticed when the model has been formulated [Aebischer:2019blw], quark and lepton sectors mutually act to prevent large deviations from SM, a welcome feature considering the small discrepancies with respect to SM emerged by measurements of flavour observables. Moreover, the possibility of having lepton flavour violation implies that modes such as those induced by $b\to s\ell_{1}^{-}\ell_{2}^{+}$ with $\ell_{1}\neq\ell_{2}$, can occur and are correlated both to purely leptonic LFV processes, $\mu^{-}\to e^{-}\gamma$, $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, $\mu^{-}\to e^{-}e^{+}e^{-}$ and $\mu^{-}\to e^{-}$ conversion in nuclei, as well as to the LFC $b\to s\ell^{-}\ell^{+}$ transitions.
To summarize our main findings, we remark that:
  • Imposing $\Delta F=2$ constraints, the NP contributions to the Wilson coefficients ${\rm Re}(C_{9}^{\rm NP})_{\ell\ell}$ and ${\rm Re}(C_{10}^{\rm NP})_{\ell\ell}$, $\ell=\mu,\,\tau$, are found to be at the $\simeq 10\%$ level of the corresponding SM coefficients. The analogous LFV coefficients ${\rm Re}(C_{9}^{\rm NP})_{\mu\tau}$ and ${\rm Re}(C_{10}^{\rm NP})_{\mu\tau}$ are of the comparable size to ${\rm Re}(C_{9(10)}^{\rm NP})_{\mu\mu}$, see Figs. 4-6.
  • The LFV $B_{(s)}$ decays are constrained by LFC modes and by purely leptonic LFV processes in a hierarchical way; the weakest bounds arise from the LFC decays, see Figs. 14-16, then by $\tau^{-}\to\mu^{-}\mu^{+}\mu^{-}$, $\mu^{-}\to e^{-}\gamma$ and by $\mu^{-}\to e^{-}e^{+}e^{-}$ and $\mu^{-}\to e^{-}$ conversion in nuclei, Figs. 17-18.
  • The role of quark and lepton observables to mutually constrain each other is visible in the results in Table 1 and in Fig. 22.
The above results are relevant for the NP searches at the new facilities planned within the future strategy for particle physics.
@@ Acknowledgements
F.D.F. thanks the coauthors of [Aebischer:2019blw]. We thank A.J. Buras for very useful discussions. This study has been carried out within the INFN project (Iniziativa Specifica) SPIF. The research has been partly funded by the European Union – Next Generation EU through the research Grant No. P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Universitá e Ricerca (MUR).
@@ Appendix A $B_{q}\to V$ local form factors
[…]
@@
Table 3: Parameters of the $B_{(s)}\to V$ form factors in Eq. (A.5) determined in [Bharucha:2015bzk].
$\langle V(p^{\prime},\epsilon)|{\bar{s}}\sigma_{\mu\nu}q^{\nu}(1+\gamma_{5})b|B_{q}(p)\rangle$ $=$ $i\epsilon_{\mu\nu\alpha\beta}\epsilon^{*\nu}p^{\alpha}p^{\prime\beta}\;2\;T_{1}(q^{2})\hskip 14.22636pt$ (A.2)
$\hskip-5.69046pt+\Big[\epsilon^{*}_{\mu}(m_{B_{q}}^{2}-m^{2}_{V})-(\epsilon^{*}\cdot q)(p+p^{\prime})_{\mu}\Big]\;T_{2}(q^{2})\hskip 19.91684pt$
$\hskip-5.69046pt+(\epsilon^{*}\cdot q)\left[q_{\mu}-\frac{q^{2}}{m_{B_{q}}^{2}-m^{2}_{V}}(p+p^{\prime})_{\mu}\right]\;T_{3}(q^{2})\,$,\,\,\,\,\,\,
@@
In our analysis we use the determination in [Bharucha:2015bzk] based on light-cone QCD sum rules (see also [Gubernari:2023puw]). Defining the ”conformal” variable
@@
with parameters $\alpha_{k}^{i}$ and $m_{R,i}$. The value of $m_{R,i}$ is close to the mass of the lightest t-channel resonance involved in the various terms in (A.1), (A.2). The parameters are determined for $V,\,A_{1},\,A_{0},\,A_{12}$ and $T_{1},\,T_{2}$, $T_{23}$ in [Bharucha:2015bzk], defining
@@
To make easier to reproduce our results, we quote the values of the parameters in Table 3, while their correlation matrix is in [Bharucha:2015bzk].
@@ Appendix B Angular coefficient functions for $B\to K^{*}(K\pi)\ell_{1}^{-}\ell_{2}^{+}$
The coefficient functions $I_{i}^{(a)}$, for $i=1,\dots 9$ and a=s,ca=s,\,c in (3.1.2), can be written in terms of the transversity amplitudes
@@
with $\lambda_{H}=\lambda(q^{2},m_{B}^{2},m_{K^{*}}^{2})$ and $\lambda_{L}=\lambda(q^{2},m_{1}^{2},m_{2}^{2})$. The expressions of the coefficient functions are:
@@ Appendix C Parametrization of the CKM and PMNS matrices
$V_{us}$ and $\gamma$ are set to $V_{us}=0.2253$ and $\gamma=64.6^{\circ}$. $|V_{cb}|$ and $|V_{ub}|$ are varied in the ranges spanned by the exclusive-inclusive measurements (73), with $|V_{cb}|_{exc}=(39.1\pm 0.5)\times 10^{-3}$, $|V_{cb}|_{inc}=(42.19\pm 0.78)\times 10^{-3}$, $|V_{ub}|_{exc}=(3.51\pm 0.12)\times 10^{-3}$, $|V_{ub}|_{inc}=(4.19\pm 0.17)\times 10^{-3}$ [HeavyFlavorAveragingGroupHFLAV:2024ctg].
@@
where $x_{ij}=\sin\theta_{ij}$ and $k_{ij}=\cos\theta_{ij}$. A global fit of the parameters in (C.3) gives values which depend on the assumed hierarchy among the neutrino masses [Esteban:2024eli]: for normal ordering
@@
The global fit in [Capozzi:2025wyn] gives consistent results. In the numerical analysis we use the parameters for normal ordering (C.4).